Mia spent one third of her money on lunch. At lunch, her friend gave her $2.50 to repay a loan. Then Mia spent $3.25 for a movie ticket and $.75 for a snack. She had $4.90 left. How much did Mia have before lunch?
Found 2 solutions by Jstrasner, mananth:
Answer by Jstrasner(112)

Hey,
For this one we need to convert this problem from words to numbers:
original money - (1/3)original money + 2.50 - 3.25 - .75 = 4.90
o.m. - (1/3)o.m. = 6.40
(2/3)o.m. = 6.40
o.m. = 6.40 x (3/2)
o.m. = 9.6
She had $9.60 before lunch

Answer by mananth(16949)

Mia spent one third of her money on lunch. At lunch, her friend gave her $2.50 to repay a loan. Then Mia spent $3.25 for a movie ticket and $.75 for a snack. She had $4.90 left. How much did Mia have before lunch?
..
let Mia have $x
she spent x/3 on lunch
so she has 2x/ 3 balance
..
friend gace 2.5 = 2 1/2 = 5/2
..
she now had 2x/3 + 5/2
she spent 3.25 = 3 1/4 = 13/4 for amovie ticket & 3/4 for a snack
2x/3 +5/2 - 13/4 -3/4
LCD = 12
(8x+30-39-9)/12 = 4.90
(8x-18)/12 =4.90
8x-18= 58.80
8x = 58.80+18
8x= 76.80
/8
x= 76.80/8
x= $9.60 amount she had
.
m.ananth@hotmail.ca